# PromptHSI: Universal Hyperspectral Image Restoration Framework for Composite Degradation

### v1

@@ 4.2 Prompt-Guided Feature Modulation
Task descriptor generation. Given a text prompt $T_{text}$ that provides the corresponding degradation description to guide the HSI restoration network to obtain the specified restoration result. In this study, we employ CLIP [40], which integrates vision and language information and performs effectively in text feature extraction. We freeze the pre-trained CLIP to maintain great linguistic consistency and add additional MLP layer $\Phi_{t}^{adpt.}$ as the feature adapter for pattern matching. With {}$\{\cdot\}_{e}$ denoting the frozen weights, this process can be expressed as:
$\footnotesize F_{text}=\Phi_{t}^{adpt.}(\{\mathcal{F}^{CLIP}_{text}\}_{e}(T_{text})),$ (6)
@@
Afterward, we design the MLP layer $\Psi_{m}^{i}$ to extract this connection with the intensity and bias embeddings for feature modulation. Therefore, it can be obtained:
$\lambda_{m}=\Psi_{m}^{I}(F_{text}),\ \mu_{m}=\Psi_{m}^{II}(F_{text}),$ (8)
where $\Psi_{m}^{I}$ and $\Psi_{m}^{II}$ are the chunk operations of $\Psi_{m}$ to form the intensity and bias controller embeddings. To address the spectral consistency challenge in HSIs, our frequency-aware feature modulation adopts an affine transformation model tailored for spectral data, formulated as:
$\hat{F}_{f}=(1+\lambda_{m}^{l})\odot F_{f}^{l}+(1+\lambda_{m}^{h})\odot F_{f}^{h}+\mu,$ (9)
@@ 4.3 Loss Function
Given the wide range of spectra inherent in HSI, we utilize only the $\ell_{1}$ loss for training. However, this strategy may not adequately capture the variability across different bands. Consequently, the $\ell_{1}$ loss might inadvertently focus on minimizing bands with larger values, affecting the spectral integrity of specific bands.
Spectral angle mapper loss. SAM metric is also critical for material identification and less sensitive to intensity variations. The $\ell_{\text{SAM}}$ loss prioritizes angular discrepancies over magnitudes, offering a reliable metric to evaluate the spectral fidelity of the reconstructed HSI.
$\ell_{\text{SAM}}=\frac{1}{N}\sum_{n=1}^{N}\cos^{-1}\left(\frac{I_{n}^{T}I^{*}_{n}+\epsilon}{\|I_{n}\|_{2}\cdot\|I_{n}^{*}\|_{2}+\epsilon}\right),$ (10)
where $I_{n}$ and $I^{*}_{n}$ represent the $n$-th spectral vector of the ground-truth and the reconstructed HSI, respectively. The term $\epsilon$ is included to avoid divergence.
Stationary wavelet transform loss. Wavelet sub-band alignment is able to manage various aspects of information encoded in HSIs. On the other words, it facilitates more detailed supervision in the frequency-domain.
$\ell_{\text{SWT}}=\sum_{j}\lambda_{j}\big{\|}\text{SWT}(I)_{j}-\text{SWT}(I^{*})_{j}\big{\|}_{1},$ (11)
where $\lambda_{j}$ represents the count of sub-bands decomposed via stationary wavelet transformation.
Band-wise MSE loss. The band-wise disparity helps stabilize the network optimization process and is advantageous to achieve convergence, balancing restoration quality across all spectra. The Band-wise MSE loss can be simply defined as follows:
$\ell_{\text{BMSE}}=\frac{1}{N}\sum_{n=1}^{N}{\|I_{n}-I_{n}^{*}\|_{2}},$ (12)
Total loss. Ultimately, the total loss function for training the proposed PromptHSI $\ell_{\text{total}}$ can be expressed as:
$\ell_{\text{total}}=\lambda_{1}\ell_{\text{1}}+\lambda_{2}\ell_{\text{SAM}}+\lambda_{3}\ell_{\text{SWT}}+\lambda_{4}\ell_{\text{BMSE}},$ (13)
where $\lambda_{1,2,3,4}$ serve as the weighting coefficients to equilibrate $\ell_{1}$, $\ell_{\text{SAM}}$, $\ell_{\text{SWT}}$, and $\ell_{\text{BMSE}}$, set at 1, 0.1, 0.01, 0.01 throughout all experiments, respectively.
@@ 5 Experimental Results
Experiment Setup. This study employed the AVIRIS dataset provided by Lin et al. [30, 18], an extensive HSI dataset, with numerous bands spanning diverse geographical areas for experiments. It includes 2,078 HSIs, which were randomly divided into training, validation, and testing sets. The training set consisted of 1,678 images and both the validation and testing sets contained 200 images each. As illustrated in Supplementary, all clean HSI was pre-processed to synthesize the composite degraded HSIs and corresponding text descriptions.
@@ 5.1 Quantitative Results and Analysis
To evaluate the performance of PromptHSI, we compared it with eight other supervised HSI restoration methods for specific degradation, including HSIDwRD [49], SERT [25], SRDNet [33], SNLSR [19], SGNet [35], AACNet [48], SST [24], and HSDT [22] in the synthesized AVIRIS dataset with composite degradations. Because there is no AiO HSI restoration method, we compare four AiO RGB image restoration methods in the experiments: AirNet [23], PromptIR [39], HAIR [5], InstructIR [10], and OneRestore [13]. The model performance was measured using the four metrics, including PSNR, SAM, RMSE, and ERGAS. Table 1 presents the quantitative results.
@@
Specifically, OneRestore [13] incorporated task description features into its encoder block, which hindered its capacity to accurately capture the spatial-spectral correlation in HSI because of the domain gap with HSI features, resulting in decoding failures. Moreover, it lacks considerations for tasks requiring high-frequency detail reconstruction, potentially limiting its effectiveness in the spatial or spectral super-resolution task.
@@ 5.2 Ablation Studies and Analysis
Integration strategy of feature-map modulation. As shown in Table 4, our proposed frequency-aware feature modulation demonstrates superior effectiveness compared to the weighted-sum manner, where we implemented an MLP layer for dynamic weighted averaging of inputs and outputs. The intensity-only configuration outperforms the bias-only, with a notably larger advantage in the ERGAS metric, suggesting that this integration strategy effectively guides network restoration across most scenarios.
Visual-prompt or text-prompt. The type of prompting is crucial, as it determines its ability to recognize various degradations and effectively guide and control model behavior. In the first configuration, we select RGB bands from HSIs as visual prompts, encoding them through CLIP [40] and using the proposed PGFM to guide the model. Secondly, we encode both RGB bands and textual prompts to generate hybrid task descriptors for guidance. In the final configuration, we remove PGFM and use the PIM and PGM block proposed in PromptIR [39] for the generation of visual prompts and interaction of features.
The experimental results are presented in Table 4. Due to multiple or composite degradations in the spatial-domain leading to serious information loss, it became difficult to obtain compact task descriptors or valid visual prompts, ultimately resulting in diminished prompt effectiveness and compromised restoration. The use of textual features helps to circumvent these issues and achieves robust prompting guidance through semantic information.

### v3

@@ 4.2 Prompt-Guided Feature Modulation
Task descriptor generation. Given a text prompt $T_{text}$ that provides the corresponding degradation description to guide the HSI restoration network to obtain the specified restoration result. In this study, we employ CLIP [48], which integrates vision and language information and performs effectively in text feature extraction. We freeze the pre-trained CLIP text encoder $\mathcal{E}^{CLIP}_{text}$ to maintain great linguistic consistency and add additional MLP layer $\Phi_{t}^{adpt.}$ as the feature adapter for pattern matching. With {}$\{\cdot\}_{e}$ denoting the frozen weights, this process can be expressed as:
$\footnotesize F_{text}=\Phi_{t}^{adpt.}(\{\mathcal{E}^{CLIP}_{text}\}_{e}(T_{text})),$ (6)
@@
Afterward, to merge text and image features, previous work [57] proposed employing a single MLP layer, $\Psi_{m}^{i}$, to integrate them with intensity and bias embeddings. Consequently, our feature modulation can be represented as:
$\footnotesize\lambda_{m}=\Psi_{m}^{I}(F_{text}),\ \mu_{m}=\Psi_{m}^{II}(F_{text}),$ (8)
where $\Psi_{m}^{I}$ and $\Psi_{m}^{II}$ are the chunk operations of $\Psi_{m}$ to form the intensity and bias controller embeddings. To address the spectral consistency challenge in HSIs, our frequency-aware feature modulation adopts an affine transformation model (Eq. 2) tailored for spectral data, formulated as:
$\footnotesize\hat{F}_{f}=(1+\lambda_{m}^{l})\odot F_{f}^{l}+(1+\lambda_{m}^{h})\odot F_{f}^{h}+\mu,$ (9)
@@ 4.3 Loss Function
Given the wide spectral range in HSIs, using only the $\ell_{1}$ loss during training may not fully capture band variability; it tends to prioritize bands with larger values, potentially compromising spectral integrity.
Spectral Angle Mapper Loss. The SAM metric is essential for material identification and is less sensitive to intensity variations. It emphasizes angular discrepancies over magnitudes:
$\footnotesize\ell_{\text{SAM}}=\frac{1}{N}\sum_{n=1}^{N}\cos^{-1}\left(\frac{I_{n}^{T}I^{*}_{n}+\epsilon}{\|I_{n}\|_{2}\cdot\|I_{n}^{*}\|_{2}+\epsilon}\right),$ (10)
where $I_{n}$ and $I^{*}_{n}$ denote the $n$-th spectral vectors of the ground-truth and reconstructed HSIs, respectively, and $\epsilon$ prevents divergence.
Stationary Wavelet Transform Loss. SWT loss enforces frequency-domain consistency via wavelet sub-band alignment:
$\footnotesize\ell_{\text{SWT}}=\sum_{j}\lambda_{j}\big{\|}\text{SWT}(I)_{j}-\text{SWT}(I^{*})_{j}\big{\|}_{1},$ (11)
with $\lambda_{j}$ weighting the contribution of each sub-band.
Band-wise MSE Loss. BMSE loss balances reconstruction across all spectral bands to stabilize optimization:
$\footnotesize\ell_{\text{BMSE}}=\frac{1}{N}\sum_{n=1}^{N}\|I_{n}-I_{n}^{*}\|_{2}.$ (12)
Total Loss. The overall loss for training PromptHSI is defined as $\ell_{\text{total}}=\lambda_{1}\ell_{\text{1}}+\lambda_{2}\ell_{\text{SAM}}+\lambda_{3}\ell_{\text{SWT}}+\lambda_{4}\ell_{\text{BMSE}}$, where the coefficients are set to $\lambda_{1}=1$, $\lambda_{2}=0.001$, $\lambda_{3}=0.01$, and $\lambda_{4}=0.01$ throughout all experiments.
@@ 5 Experimental Results
Experiment Setup. This study employed the AVIRIS dataset provided by Lin et al. [35, 20], an extensive HSI dataset with numerous bands spanning diverse geographical areas, for experiments. The dataset comprises 2,078 HSIs, which were randomly divided into training (1,678 images), validation (200 images), and testing (200 images) sets. As illustrated in the Supplementary, all clean HSIs were pre-processed to synthesize composite degraded HSIs along with corresponding text descriptions. This dataset is publicly available to facilitate further research.
@@ 5.1 Quantitative Results and Analysis
To evaluate the performance of PromptHSI, we compared it with eight other supervised HSI restoration methods for specific degradation, including HSIDwRD [60], SERT [29], SRDNet [39], SNLSR [21], SGNet [43], AACNet [56], SST [28], and HSDT [26] in the synthesized AVIRIS dataset with composite degradations. Because there is no AiO HSI restoration method, we compare four AiO RGB image restoration methods in the experiments: AirNet [27], PromptIR [47], HAIR [4], InstructIR [10], AdaIR [11], and OneRestore [15]. The model performance was measured using the four metrics, including PSNR, SAM, RMSE, and ERGAS. Table 1 presents the quantitative results.
@@
Specifically, OneRestore [15] integrated task description features into its encoder, which hindered its ability to capture spatial-spectral correlations in HSIs due to a domain gap, leading to decoding failures.
@@ 5.2 Ablation Studies and Analysis
Integration strategy of frequency-aware modulation. As shown in Table 12, modulating frequency via prompting demonstrates superior effectiveness compared to the weighted-sum manner, where we implemented an MLP layer for dynamic weighted averaging of inputs and outputs. The intensity-only configuration outperforms the bias-only, with a notably larger advantage in the ERGAS metric, suggesting that this integration strategy effectively guides network restoration across most scenarios.
Visual-prompt or text-prompt. The type of prompting is crucial for recognizing degradations and guiding model behavior. In our first configuration, we use RGB bands from HSIs as visual prompts, encoded via CLIP [48] and guided by PGFM. In the second, we combine RGB and textual prompts to form hybrid task descriptors. In the final configuration, PGFM is removed and we employ the PIM and PGM blocks from PromptIR [47] for generating visual prompts and feature interaction.
Table 4 shows that composite spatial degradations cause significant information loss, making it difficult to obtain compact task descriptors or effective visual prompts. Incorporating textual features overcomes these challenges by providing robust semantic guidance.